OpenStudy (anonymous):

can you reduce 1.5/100?

OpenStudy (anonymous):

15/1000

OpenStudy (anonymous):

yea we can reduce it , 1.5/100 = 15/1000

OpenStudy (anonymous):

1.5/100=15/1000=3/200

OpenStudy (anonymous):

i mean, it would just either move the decimal or become 3/200